Science Report – May 10th

Crew 212-LATAM-III

Submitted by Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 5

Sim suit used for first EVA: 7, 4 and 10.
used for 1 hour

Sim suit: 4
Voltage level before EVA: 12.6 V
Voltage level after EVA: 8.4 V

Sim suit: 7
Voltage level before EVA: 13.1 V
Voltage level after EVA: 12.7 V

Sim suit: 10
Voltage level before EVA: 13.1 V
Voltage level after EVA: 12.9 V

Sim suit used for second EVA: 3, 5, 8 and 9.
used for 1 1/2 hours

Sim suit: 3
Voltage level before EVA: 13.4 V
Voltage level after EVA: 12.8 V

Sim suit: 5
Voltage level before EVA: 13.4 V
Voltage level after EVA: 12.8 V

Sim suit: 8
Voltage level before EVA: 13.5 V
Voltage level after EVA: 12.3 V

Sim suit: 9
Voltage level before EVA: 13.3 V
Voltage level after EVA: 12.8 V

General comments:
– It seems that the battery of sim suit 4 is draining the battery very fast.
– The battery recharger was found. I will proceed to charge the extra batteries and see if any of them have gone bad

Operations Report – May 10th

Crew 212 Operations Report 10-MAY-2019

SOL: 5

Name of person filing report: Paolo Guardabasso

Non-nominal systems: None

Notes on non-nominal systems: Nothing to report

Generator (hours run last night): [From Director]

On (last night): 9:00 PM

Off (this morning): [From Director]

Solar— SOC 99% at 7:40 PM

Diesel Reading – % [From Director]

Propane Reading – % [From Director]

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 200 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 155 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143454.5

Toilet tank emptied: No

Deimos rover used: no (out of order)

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes

Hours: 92.6h

Beginning charge: 100%

Ending charge: 57%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 59.1h

Beginning charge: 100%

Ending charge: 74%

Currently charging: No

Curiosity rover used: Yes

Hours: 94.7h

Beginning charge: 100%

Ending charge: 75%

Currently charging: Yes

Notes on rovers: Spirit showed an anomalous power consumption, hitting 70% charge after about 10 minutes of driving. It forced the EVA crew to come back to the Hab, where the charge was 57%
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report

CrewCar used and why, where? Off Campus

General notes and comments: We received the visit from two journalists, who toured the station and interviewed the crew members. One EVA was performed this morning and the crew was accompanied by two journalists. The EVA in the afternoon postponed because of bad weather, but it was successfully carried out 30 minutes later.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: It was depressurized to allow the EVA crew to take the drones for the experiments.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: a crew member had a minor injury.

Questions, concerns, and requests to Mission Support: Following the issue of the heater, we were aware of the procedure in the handbook and we already tried keeping the hot water sink fully on for more than 10 seconds, but the heater still didn’t activate. Both gas and water are at 100%, and the propane and water lines valves are open.

EVA Report – May 10th

Crew 212 EVA Report 10-MAY-2019

EVA #06

Author: Zoe Townsend

Purpose of EVA: First testing of drones and soils collection

Start time: 15:00

End time: 17:00


Today’s EVA was to run two simultaneous experiments. The first experiment was to collect sand and rocks from various points of interest around the Hab. This is combined with the soil (red clay) collected in EVA #4. This soil, sand and rock solution will be mixed with the bacteria from radioactive regions to help with plant growth.

The second experiment was the continuation of the Drone testing, The X-5 and Tarot were tested today, the Tarot was flown around the Hab at different points that were identified as points of interest in EVA #4. Cameras on board took videos of the flight until battery issues prevented further flying. Secondly, the X-5 was flown in an empty area north of the observatory, the flight was sustained but cut short after a maneuver caused an unplanned landing.

Notes from the debriefing:

· Bring water

· Ensure that all the roles are fully defined prior to the start of the EVA

· Standardize the depressurizing time of the EVA (3 mins)

· Ensuring that there is confirmation from the Hab prior to search and recover of Drone.

Coordinates (UTM NAD27 CONUS): Sector12S Northern 4250933, Eastern 518167

EVA Participants: Hector, Vittorio (Medic), Paolo and Zoe (Commander)

Road(s) and routes per MDRS Map: 1103

Mode of travel: (walking or driving)? Walking

EVA Report – May 10th

Crew 212 EVA Report 10-MAY-2019

EVA #05

Author: Marlen Castillo Vilcahuaman

Purpose of EVA: Locating the Martian flora according to Paul Sokoloff’s study, to obtain their soil for the study of microbiological diversity in desert soils. We’ll take photos of the flowers, characterize and pick up their soil.

Start time: 10:30 AM
End time: 11:30 AM


The purpose of this EVA was to find the Martian flora registered in Sokoloff’s paper. Although we have no work in plant ecology for this mission, we thought soil samples from flowers would have an interesting diversity in microbes, which would mean more morphological diversity in the petri dish. We asked for two rovers (Spirit and Curiosity).

However, the LA Times visit altered our plans. We took Opportunity too so that they could come with us and watch our fieldwork during EVAs. Our final point was Sector12S Northern 4250933, Eastern 518167. But we had problems with Spirit. Just after turning in Galileo Road 1104, one of our crew members reported that Spirit was with 70% of battery.

We were surprised by this battery usage. Hence, we decided to stop at that point (12S 518779.59 m E 4251383.04 m N) and do our recollection work, collecting soils and flowers and doing some photography of them. After 20 minutes of walking and going around the field in search of flowers, we decided to turn around and go back to the Hab.

Spirit arrived at the Hab with 58% of battery. Curiosity’s return was nicer, with its battery being higher than 60%.

We collected four soil samples, identifying them per plant.

Destination: URC North Site and continuing in the first part of the road on the east.

Coordinates (UTM NAD27 CONUS):

Looked forward to Sector12S Northern 4250933, Eastern 518167
Stayed in: 12S 518779.59 m E 4251383.04 m N

EVA Participants: Marlen (Commander), Mariona (EVA’s HSO), Camilo

Road(s) and routes per MDRS Map: 1103, Galileo Road 1104

Mode of travel: driving and walking

Greenhab Report – May 10th

Crew 212 GreenHab Report – 10-May-2019: SOL 5

GreenHab Officer: Hector Palomeque

Environmental control: Ambient with the door open from 07:00 to 14:30 + door closed from 14:30 to 15:30 + door open from 15:30 to the rest of the day

Shade cloth 80%: on

Average temperatures:
Low: 15.0 °C (±0.85)
High: 32.7 °C (±1.70)

Hours of supplemental light: 0

Daily water usage for crops: 5 gallons

Water in Blue Tank: 157 gallons

Time(s) of watering for crops: 12:00

Changes to crops: All plants look healthy. Half of Basil plants have already bloomed. Tomatoes in the wood table (black pots) still blooming. Tomatoes in the blue tray have already started fructification (1 tomato).

Narrative: The GreenHab door had to been closed at 14:30 due to the rain and then re-opened at 15:30 because the weather has changed dramatically. Plants were watered just once at noon. The usual watering at early morning (at 07:00) was avoided due to the substrate has remained moist since the day before.

Some herbs and Chard were harvested for preparing the lunch of the Crew.

Harvested: 100g Chard + 5g Parsley + 6g Basil

Support/supplies needed: None.


[status draft

Good evening,

Today was Sol 5 and we had two visitors from the LA Times. The day started with the meet and greet followed by a tour of the station.

An attempt was made to use the solar telescope, but the clouds prevented observation.

Today two EVA’s took places. The first was a midmorning one to explore 3 sights on Galileo road, the URC North sight. This was completed with three members of the crew and shadowed by the two reporters.

The aim was to find plants to identify according to Paul Sokoloff study, however, problems with the Spirit rover meant that the charge ran down to 60% after the first visit and therefore the other two places will require a further EVA.

After the first EVA, we prepared lunch using cultivated foods from the GreenHab, the Green and Swiss chard was mixed into a tomato sauce and infused with Basil and Parsley also from the GreenHab. This was served to our guests alongside the leftover desserts from the Latin American cultural night.

The second EVA of the day was to run two simultaneous experiments, the first was to collect soil and rock samples to mix with previously collected soil. The mixture would then be combined with the radiated bacteria for one of the experiments we are running. The other experiment was a continuation from EVA#4, where the X-5 drone and the Tarot were flown with cameras around the Hab exterior.

Once we returned we completed another round of the problem-solving game.

Kind Regards,
Zoe Townsend

Sol Summary – May 10th

Crew 212 Sol 5 Summary Report 10-MAY-2019

Sol: 5

Summary Title: After the rain, it’s always the sun again

Author’s name: Marlen Castillo Vilcahuaman

Mission Status: A-ok

Sol Activity Summary: There were 2 EVAs programmed for today. We had some issues with Spirit’s battery (see Operations report) that’ll we’ll need to keep an eye on. We’ll see how this rover handles itself in the following days. Oh. And we also had the visit of the LA Times!

Look Ahead Plan: 1 EVA for tomorrow, and new creative ways to prepare our meals (we discovered that this is an entertaining task!)

Anomalies in work: None

Weather: We had a cloudy day. After some heavy rain, now the sun is sure shining!

Crew Physical Status: A crewmember had an incident (please see HSO report)

EVA: 1

Reports to be file: Operations Report, GreenHab Report, Journalist Report, Daily Photo Report, EVA Request, Space Suit Report, EVA Report.

Support Requested: The Non-Burnable trash bag needs to be picked up. We’re also looking for some greens in our food, if possible. Broccoli is running out.

Copyright © The Mars Society. All rights reserved. | Main Site